OpenStudy (anonymous):

given sec theta = sqrt 10 and tan theta =3, determine sin (90 deg-theta)

OpenStudy (loser66):

sin (90 - theta) = cos theta but I don't know why they give you many information like that, because sec = 1/ cos, \[\frac {1}{cos\theta}=\sqrt 10\]therefore \[cos\theta= \frac{1}{\sqrt10}\]and nothing to do with tan theta. wonder why and how?

OpenStudy (anonymous):

me to im tryna understand it also

OpenStudy (loser66):

is there any question else? relate to this problem?

OpenStudy (anonymous):

is this undefined u think

OpenStudy (anonymous):

no its a past paper practice quection im practicng for an exam

OpenStudy (anonymous):

one of the multiple choice has undefined as one of the answers

OpenStudy (loser66):

it's not undefined.

OpenStudy (anonymous):

ohhh

OpenStudy (loser66):

figure out something? I doubt myself

OpenStudy (anonymous):



OpenStudy (loser66):

90 is pi/2 not pi friend

OpenStudy (anonymous):

oops my bad

OpenStudy (anonymous):

read it wrong